OpenStudy (anonymous):

A spherical balloon is inflated with helium at the rate of 100pie ft^3/min/ a) how fast is the balloon's radius increasing at the interest the radius is 5 feet? b) how fast is the surface area increasing at that instant.

OpenStudy (anonymous):

Given dV/dt = 100pi Volume of sphere, V= 4/3 pi r^3 dV/dr = 4 pi r^2 To get the rate of increase of the radius when radius is 5 ft, note that, dr/dt = (dV/dt) * (dr/dV) multiplying it like this, the dV gets cancelled, giving you the dr/dt which is what you want. dr/dt = 100pi / 4pi r^2 = 25/ r^2 Substitute r= 5ft, dr/dt = 25/ 25= 1 ft/min Rate of increase of the surface area when radius is 5ft, Surface area, A =4pi r^2 dA/dr = 8pi r Same as the above, To get dA/dt = dA/dr * dr/dt dA/dt = 8 pi r * 1 = 8 pi r Substitute r = 5 ft. dA/dt = 40pi ft^2 /min

OpenStudy (anonymous):

@tiaph can you help me on one like this